Supplemental Operations Report – May 6th

Supplemental Operations Report 06May2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation.

Power system:
Solar: nominal. Input breaker of one of MPTT was tripped several times this week. After resetting the breaker and solar charger controller current work is nominal.SOC Last 24 hours: Max 100%; Min 65 %; Avg 77.3%.
VDC Last 24 hours: Max 58.99V; Min 45.69; Avg 51.66V.Generator run time: about 12am – 6am on May 4-5th. Generator hours 3787.4.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 60%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 76%

Water (Static Tank) – 289 gallons
Water in GreenHab – 99 gallons
Water (Outpost tank) – 275 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: No.
Hours: 191.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for water runs and supplies pick up.

Crew Car used and why, where? No. In Grand Junction.

General notes and comments: Nominal.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Headset #3 was replaced.

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Nominal.

Summary of GreenHab Operations: Nominal, heater turned off.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation .

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Director’s trailer – small leak was found in hot and cold water line of the bathroom sink. It is very minor and in the L shape connection right under the floor, it can be seen from the storage compartment. It does not require immediate actions but should be taken care of in the future.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site